The volume, V cm³, of a metallic cube of side length x cm, is increasing at the constant rate of 0.216 cm3 s^-1 .


a)Determine the rate at which the side of the cube is increasing when the side length reaches 6 cm.


b) Find the rate at which the surface area of the cube, A cm², is increasing when the side length reaches 6 cm.


1
Expert's answer
2022-03-07T12:04:03-0500

(a)


"V=x^3"

Differentiate both sides with respect to "t"


"\dfrac{dV}{dt}=3x^2\dfrac{dx}{dt}"

Then


"\dfrac{dx}{dt}=\dfrac{1}{3x^2}\dfrac{dV}{dt}"

Given "\dfrac{dV}{dt}=0.216\ cm^3\cdot s^{-1}, x=6\ cm"


"\dfrac{dx}{dt}=\dfrac{1}{3(6\ cm)^2}(0.216\ cm^3\cdot s^{-1})=0.002\ cm\cdot s^{-1}"

The side of the cube is increasing at the of "0.002\ cm\cdot s^{-1}."


(b)


"A=x^2"

Differentiate both sides with respect to "t"


"\dfrac{dA}{dt}=2x\dfrac{dx}{dt}"

Given "\dfrac{dx}{dt}=0.002\ cm\cdot s^{-1}, x=6\ cm"


"\dfrac{dA}{dt}=2(6\ cm)(0.002\ cm\cdot s^{-1})=0.024\ cm^2\cdot s^{-1}"

The surface area of the cube is increasing at the of "0.024\ cm^2\cdot s^{-1}."